A distribution of scores on a math exam has a mean of 88 and a standard deviation of 12. The instructor would like to curve the exam by adding 2 points to all exams. What will the new mean, variance, and standard deviation be
Answer #1

New mean = old mean + added value

where old mean = 88 and added value = 2

new mean = 88+2 = 90

standard deviation will remain same, i.e. 12 because the difference between the data values is unchanged.

and variance = (standard deviation)^2 = 12^2 = 144, i.e. it is also same as earlier as the difference between data values is unchanged.